Calculating returns and standard deviation. Based on the following information, can you calculate the expected return and standard deviation for the two stocks?:

State of economy. Prob of st of econ Rate of return if state occurs

  Stock A Stock B

Recession    .25 .06 -.20    

Normal .55 .07 .13

Boom .20 . .11 .33
Answer #1

The values provided for stock A and stock B in the question are putting in the table as follows:

FOR STOCK A
STATE OF ECONOMY PROBABILITY RETURN PROBABILITY * RETURN (RETURN-EXPECTED RETURN) SQUARED OF (RETURN -EXPECTED RETURN) PROBABILITY * SQUARED OF (RETURN -EXPECTED RETURN)
A B C D=B * C E=C-D F=E^2 G= B * F
Recession 0.25 6 1.5 4.5 20.25 5.0625
Normal 0.55 7 3.85 3.15 9.9225 5.457375
Boom 0.2 11 2.2 8.8 77.44 15.488
0
TOTAL 1 EXPECTED RETURN OF STOCK A =7.55 0
VARIANCE=26.007875
FOR STOCK B
STATE OF ECONOMY PROBABILITY RETURN PROBABILITY * RETURN (RETURN-EXPECTED RETURN) SQUARED OF (RETURN -EXPECTED RETURN) PROBABILITY * SQUARED OF (RETURN -EXPECTED RETURN)
A B C D=B * C E=C-D F=E^2 G= B * F
Recession 0.25 -20 -5 -15 225 56.25
Normal 0.55 13 7.15 5.85 34.2225 18.822375
Boom 0.2 33 6.6 26.4 696.96 139.392
0
TOTAL 1 EXPECTED RETURN OF STOCK B =8.75 0
Variance of stock A=26.007875
Variance of stock B=214.464375
standard deviation for the stocks B = Square Root of variance of stock B=14.64%
VARIANCE=214.464375
standard deviation for the stocks A = Square Root of variance of stock A =5.10%